OpenStudy (anonymous):

What is the midpoint of the segment with the endpoints (–5,2) and (–3,–4)?

OpenStudy (anonymous):

-4,-1

OpenStudy (amistre64):

its the avergare of the xs and ys

OpenStudy (amistre64):

average ....

OpenStudy (anonymous):

x =(x1 +x2)/2 and y=(y1+y2)/2